OpenStudy (anonymous):

Solve M = 2x + 3y for y

Parth (parthkohli):

1) Subtract \(2x\) from both sides. 2) Divide both sides by \(3\).

OpenStudy (anonymous):

so its this right? \[\frac{\cancel{3}y}{\cancel{3}}=\frac{M-2x}{3}\]

Parth (parthkohli):

That's correct!

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20110208103428AAPbtD3 here is how you can solve it

OpenStudy (anonymous):

M=2x+3y Your goal should be to isolate y on one side M - 2x = 2x + 3y -2x Subtract 2x from both sides of the equation M - 2x = 3y (M-2x)/3=3y/3 Divide both sides of the equation by 3 (M-2x)/3=y y=M/3-2x/3 or y=(M-2x)/3

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

wellcome :) anytime you need help message me because i am doing the same lessons

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

ugotnukd you are welcome...

Parth (parthkohli):

G'job folks!